OpenStudy (anonymous):

Find the value of n + 6 how would you do this one?

OpenStudy (anonymous):

depends on n...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hexagon - so summ of the internal angles = 720 degrees form an equation in n and solve

OpenStudy (anonymous):

so its n + n + 6 +.......... = 720

OpenStudy (paxpolaris):

yes

OpenStudy (paxpolaris):

sum of angles of a polygon with x number of sides: 180*(x-2) degrees

OpenStudy (karatechopper):

488+n+n+6 488+2n+6 494+2n=720 494-494+2n=720-494 2n=226 2n/2=226/2 n=113

OpenStudy (karatechopper):

is tht right?

OpenStudy (anonymous):

yea i think so

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

n+ 6 = 119